# Classification of coadjoint orbits for the maximal unipotent subgroup in the simple group of type $F_{4}$

### v4

@@ 3. Classification of orbits
$\exp(x).\lambda=\sum\limits_{\alpha\in\Phi^{+}}\lambda\left(\sum\limits_{i=0}^{\infty}\dfrac{(-1)^{i}}{i!}\underbrace{[x,[x,\dots,[x,e_{\alpha}]\dots]]}_{\text{$i$ commutators}}\right)\otimes e_{\alpha}^{*}.$
Denote by $\chi$ the element $\chi=\sum\limits_{\alpha\in\Phi^{+}}x_{\alpha}\otimes e_{\alpha}\in M$. For a form $\lambda\in\mathfrak{n}^{*}$, we will consider the element $\exp(\chi).\lambda\in M^{*}$. Let us also set $\overline{\mathfrak{n}}=\overline{\mathbb{F}}\otimes_{\mathbb{F}}\mathfrak{n}$, $\overline{N}=\overline{\mathbb{F}}\otimes_{\mathbb{F}}N$. Note that if $\widetilde{x}=\sum\limits_{\alpha\in\Phi^{+}}\widetilde{x}_{\alpha}e_{\alpha}\in\overline{\mathfrak{n}}$ with $\widetilde{x}_{\alpha}\in\overline{\mathbb{F}}$, then $\exp(\widetilde{x}).\lambda=(\exp(\chi).\lambda)(\widetilde{x}_{\alpha},\alpha\in\Phi^{+})$.
@@
Theorem 3.2. Let $\lambda\in S$, $D=\textup{supp}(\lambda)$, and let $\gamma\in\Phi^{+}$ satisfy condition (1). Then there exists a rational function $F_{D,\gamma}\in\overline{\mathbb{F}}(y_{\beta},\beta\succ\gamma)$ such that
@@
where $\mu=\exp(\chi).\lambda\in M^{*}$.
@@
Example 3.3. Consider the same support as in Example Classification of coadjoint orbits for the maximal unipotent subgroup in the simple group of type $F_{4}$. Let $D=\textup{supp}(\lambda)=\{\alpha_{4},\alpha_{3}+\alpha_{4},\alpha_{2},\alpha_{2}+2\alpha_{3}\}$. We will show that for this $\lambda$ and for all roots $\gamma$ satisfying condition (1), formula (2) holds. Let $\mu=\exp(\chi).\lambda,~\lambda=\sum\limits_{\alpha\in\Phi^{+}}\lambda_{\alpha}e_{\alpha}^{*},~\mu=\sum\limits_{\alpha\in\Phi^{+}}\mu_{\alpha}\otimes e_{\alpha}^{*}$. Let us write down explicitly the coordinates of interest. As in Example Classification of coadjoint orbits for the maximal unipotent subgroup in the simple group of type $F_{4}$, we denote the structural constants in the relation $Ce_{\alpha+\beta}=[e_{\alpha},e_{\beta}]$ by $C_{\alpha,\beta}$. Their values can be found in [KP23]. We have
$\begin{split}\mu_{\alpha_{2}+2\alpha_{3}}&=\lambda(e_{\alpha_{2}+2\alpha_{3}})=\lambda_{\alpha_{2}+2\alpha_{3}};\\ \mu_{\alpha_{2}+\alpha_{3}+\alpha_{4}}&=0;\\ \mu_{\alpha_{2}+\alpha_{3}}&=-\lambda([x,e_{\alpha_{2}+\alpha_{3}}])=-\lambda([x_{\alpha_{3}}e_{\alpha_{3}},e_{\alpha_{2}+\alpha_{3}}])=-C_{\alpha_{3},\alpha_{2}+\alpha_{3}}x_{\alpha_{3}}\lambda(e_{\alpha_{2}+2\alpha_{3}})=2x_{\alpha_{3}}\lambda_{\alpha_{2}+2\alpha_{3}};\\ \mu_{\alpha_{2}}&=\lambda(e_{\alpha_{2}})+\dfrac{1}{2}\lambda([x,[x,e_{\alpha_{2}}]])=\lambda_{\alpha_{2}}+\dfrac{1}{2}\lambda([x_{\alpha_{3}}e_{\alpha_{3}},[x_{\alpha_{3}}e_{\alpha_{3}},e_{\alpha_{2}}]])\\ &=\lambda_{\alpha_{2}}+\dfrac{1}{2}C_{\alpha_{3},\alpha_{2}+\alpha_{3}}C_{\alpha_{3},\alpha_{2}}x_{\alpha_{3}}^{2}\lambda(e_{\alpha_{2}+2\alpha_{3}})=\lambda_{\alpha_{2}}-x_{\alpha_{3}}^{2}\lambda_{\alpha_{2}+2\alpha_{3}}=\lambda_{\alpha_{2}}-\dfrac{\mu_{\alpha_{2}+2\alpha_{3}}^{2}}{4\mu_{\alpha_{2}+2\alpha_{3}}};\\ \mu_{\alpha_{3}+\alpha_{4}}&=\lambda(e_{\alpha_{3}+\alpha_{4}})=\lambda_{\alpha_{3}+\alpha_{4}};\\ \mu_{\alpha_{4}}&=\lambda(e_{\alpha_{4}})-\lambda([x,e_{\alpha_{3}+\alpha_{4}}])=\lambda_{\alpha_{4}}-\lambda([x_{\alpha_{3}}e_{\alpha_{3}},e_{\alpha_{3}+\alpha_{4}}])\\ &=\lambda_{\alpha_{4}}-C_{\alpha_{3},\alpha_{4}}x_{\alpha_{3}}\lambda(e_{\alpha_{3}+\alpha_{4}})=\lambda_{\alpha_{4}}+x_{\alpha_{3}}\lambda_{\alpha_{3}+\alpha_{4}}=\lambda_{\alpha_{4}}+\dfrac{\mu_{\alpha_{2}+\alpha_{3}}\mu_{\alpha_{3}+\alpha_{4}}}{2\mu_{\alpha_{2}+2\alpha_{3}}}.\end{split}$
@@
In fact, the validity of formula (2) in this example could have been understood without writing out the coordinates explicitly. It is enough to examine Table 1. Since $\lambda(e_{\alpha_{2}+2\alpha_{3}})=\mu(e_{\alpha_{2}+2\alpha_{3}})$ and $\mu(e_{\alpha_{2}+\alpha_{3}})$ involves only $x_{\alpha_{3}}$ and $\lambda(e_{\alpha_{2}+2\alpha_{3}})$, the variable $x_{\alpha_{3}}$ can be expressed as a rational function of $\mu(e_{\alpha_{2}+\alpha_{3}})$ and $\mu(e_{\alpha_{2}+2\alpha_{3}})$. Consequently, since $\mu(e_{\alpha_{2}})-\lambda(e_{\alpha_{2}})$ depends only on $x_{\alpha_{3}}$ and $\lambda(e_{\alpha_{2}+2\alpha_{3}})$, it can be expressed as a rational function of $\mu(e_{\alpha_{2}+\alpha_{3}})$ and $\mu(e_{\alpha_{2}+2\alpha_{3}})$. Furthermore, $\mu(e_{\alpha_{3}+\alpha_{4}})=\lambda(e_{\alpha_{3}+\alpha_{4}})$. As $\mu(e_{\alpha_{4}})-\lambda(e_{\alpha_{4}})$ involves only $x_{\alpha_{3}}$ and $\lambda(e_{\alpha_{3}+\alpha_{4}})$, it can be expressed as a rational function of $\mu(e_{\alpha_{2}+\alpha_{3}})$ and $\mu(e_{\alpha_{3}+\alpha_{4}})$.
@@
Now let us proceed to the proof of Theorem Classification of coadjoint orbits for the maximal unipotent subgroup in the simple group of type $F_{4}$. First, we prove it in the special case of forms whose support consists of a single root. Any such form obviously belongs to $S$.
Lemma 3.4. Theorem Classification of coadjoint orbits for the maximal unipotent subgroup in the simple group of type $F_{4}$ holds for any $\lambda\in\mathfrak{n}^{*}$ with $|\textup{supp}(\lambda)|=1$.
@@
Proof. Denote $D=\textup{supp}(\lambda)=\{\delta\}$. Note that in this case, condition (1) depends only on $D$, not on the specific $\lambda$, and it holds for those $\gamma$ that are not singular to $\delta$. Let $\mu=\exp(\chi).\lambda$. We need to prove that for any $\gamma\notin S_{\delta}$, the coordinate $\mu(e_{\gamma})$ is a function of the coordinates $\mu(e_{\beta}),\beta\succ\gamma$}. Observe that if $\gamma\succ\delta$, then $\mu(e_{\gamma})=0$, and if $\gamma=\delta$, then $\mu(e_{\delta})=\lambda(e_{\delta})$. For these two cases, the condition is obviously satisfied. It remains to prove it for $\gamma\prec\delta$.
@@
The coordinate $\mu(e_{\gamma})$ is a polynomial in the $x_{\beta}$ with coefficients from $\overline{\mathbb{F}}$. It does not contain the variable $x_{\beta_{2k-i+1}}$ if $\gamma\succ\beta_{i}$. Let $\beta_{2k-j+1}$ be the largest root in $S_{\delta}$ such that $x_{\beta_{2k-j+1}}$ appears in $\mu(e_{\gamma})$. Using formula (3), we express $x_{\beta_{2k-j+1}}$ in terms of $\mu(e_{\beta_{j}})$ and $x_{\alpha},\alpha<\beta_{2k-j+1}$}. Thus, $\mu(e_{\gamma})$ becomes a polynomial that no longer involves $x_{\beta_{2k-j+1}}$. We perform a similar elimination successively for $\beta_{2k-j+2},\beta_{2k-j+3}$, and so on, until $\beta_{2k}$. Note that at each step, a variable $x_{\beta_{i}}$ that was initially absent or already removed cannot reappear, because we proceed in decreasing order of the singular roots. As a result, $\mu(e_{\gamma})$ is represented as a polynomial that does not contain $x_{\beta},\beta\in S_{\delta}$, i.e., $\mu(e_{\gamma})=G(\mu(e_{\beta_{i}}),i=1,\ldots,j;~x_{\alpha},\alpha\notin S_{\delta})$ for some polynomial $G\in\overline{\mathbb{F}}[t_{\beta_{i}},i=1,\ldots,j;~x_{\alpha},\alpha\notin S_{\delta}]$. Observe that $\beta_{i}\succ\gamma$ for all $i=1,\ldots,j$.
@@
By our assumption, $\mu(e_{\gamma})$ cannot be expressed solely in terms of the subsequent coordinates. This implies that $G$ is not identically constant and actually depends on at least one variable $x_{\alpha}$. Then there exists a non-empty Zariski-open subset $U\subseteq\overline{\mathbb{F}}^{j}$ such that for any $c=(c_{1},c_{2},\ldots,c_{j})\in U$, the specialized polynomial $G(c_{i},i=1,\ldots,j;~x_{\alpha},\alpha\notin S_{\delta})$ is non-constant. Fix such a tuple $c_{1},c_{2},\ldots,c_{j}\in\overline{\mathbb{F}}$. Choose an index $\varepsilon\notin S_{\delta}$ and constants $c_{\alpha}\in\overline{\mathbb{F}},\alpha\notin S_{\delta},\alpha\neq\varepsilon$, so that the one-variable polynomial $g(x_{\varepsilon})=G(c_{i},i=1,\ldots,j;~x_{\varepsilon},c_{\alpha},\alpha\notin S_{\delta},\alpha\neq\varepsilon)$ is non-constant. The field is algebraically closed, for any chosen $c_{0}\in\overline{\mathbb{F}}$ there exists a $c_{\varepsilon}\in\overline{\mathbb{F}}$ such that $g(c_{\varepsilon})=c_{0}$. Fix $c_{0}$ and the corresponding $c_{\varepsilon}$. Now, choose arbitrary values $c_{j+1},\ldots,c_{2k}\in\overline{\mathbb{F}}$. Substituting the values $c_{\varepsilon},c_{\alpha},\alpha\notin S_{\delta}$ into formula (3), we can determine values for $x_{\beta_{2k}},x_{\beta_{2k-1}},\ldots,x_{\beta_{1}}$ sequentially, in such a way that $\mu(e_{\beta_{1}})=c_{1},\mu(e_{\beta_{2}})=c_{2},\ldots,\mu(e_{\beta_{2k}})=c_{2k}$. For this assignment, we obtain $\mu(e_{\gamma})=G(c_{i},i=1,\ldots,j;~c_{\varepsilon},c_{\alpha},\alpha\notin S_{\delta})=c_{0}$.
@@
Therefore, there exists a non-empty Zariski-open subset $V\subseteq\overline{\mathbb{F}}^{2k+1}$ such that for any $(c_{0},c_{1},\ldots,c_{2k})\in V$ , we can find an element $\overline{\mu}\in\overline{N}.\lambda$ whose coordinates $\overline{\mu}(e_{\gamma}),\overline{\mu}(e_{\beta_{1}}),\ldots,\overline{\mu}(e_{\beta_{2k}})$ take these values. The orbit $\overline{N}.\lambda$ is an algebraic variety. Our construction yields a surjection from $\overline{N}.\lambda$ to $\overline{\mathbb{F}}^{2k+1}$, which implies $\dim\overline{N}.\lambda\geq 2k+1$. However, $\dim\overline{N}.\lambda=\dim\overline{\mathfrak{n}}.\lambda=\dim\mathfrak{n}.\lambda=|S_{\delta}|=2k$. This contradiction completes the proof. $\square$
We introduce further notation. Let $\alpha,\beta\in\Phi^{+}$. Denote
@@
$T_{\alpha,\beta}(x):=(\exp(\chi).e_{\beta}^{*})(e_{\alpha})\in\overline{\mathbb{F}}[x_{\gamma},\gamma\in\Phi^{+}]$. It is clear that if $\lambda\in\mathfrak{n}^{*},\alpha\in\Phi^{+}$, then
@@
$\begin{split}(\exp(\chi).\lambda)(e_{\alpha})=\left(\exp(\chi).\left(\sum\limits_{\beta\in\textup{supp}(\lambda)}\lambda(e_{\beta})e_{\beta}^{*}\right)\right)(e_{\alpha})&=\sum\limits_{\beta\in\textup{supp}(\lambda)}\lambda(e_{\beta})(\exp(\chi).e_{\beta}^{*})(e_{\alpha})\\ &=\sum\limits_{\beta\in\textup{supp}(\lambda)}\lambda(e_{\beta})T_{\alpha,\beta}(x).\end{split}$
@@
Now consider the following construction (cf. the construction of the set $\zeta_{D,\gamma}$ used in the proof of the sufficient condition in Proposition Classification of coadjoint orbits for the maximal unipotent subgroup in the simple group of type $F_{4}$). Let $\lambda\in S,~D=\textup{supp}(\lambda)$, and let the root $\gamma\in\Phi^{+}$ satisfy condition (1). Starting from $D$ and $\gamma$, we construct a set $\zeta^{\prime}_{D,\gamma}\subset\Phi^{+}\times\Phi^{+}$: by definition, it consists of those pairs $(\alpha,\beta)$ for which the polynomial $T_{\alpha,\beta}(x)$ can be expressed as a rational function of $\{(\exp(\chi).\lambda)(e_{\delta})\}_{\delta\succ\gamma}$.
@@
Example 3.5. Table 1 helps visualize the structure of the polynomials $T_{\alpha,\beta}(x)$. One simply examines the cell in row $\alpha$ and column $\beta$. For each pair $\gamma,\delta\in\Phi^{+}$ such that $\gamma+\delta\in\Phi^{+}$, denote the corresponding non-zero structure constant by $C_{\gamma,\delta}$, so that $e_{\gamma},e_{\delta}]=C_{\gamma,\delta}e_{\gamma+\delta}$. Each monomial in $T_{\alpha,\beta}(x)$ has the form
@@
where $\gamma_{1},\ldots,\gamma_{n}\in\Phi^{+},~\alpha+\gamma_{1}+\ldots+\gamma_{i}\in\Phi^{+}$ for all $i$ from 11 to $n$, and $\alpha+\gamma_{1}+\ldots+\gamma_{n}=\beta$. Hence, it suffices to consider all possible decompositions of $\beta-\alpha$ into a sum of positive roots. For example, $T_{\alpha_{3},\alpha_{2}+2\alpha_{3}+\alpha_{4}}(x)=-x_{\alpha_{2}+\alpha_{3}+\alpha_{4}}+\dfrac{1}{2}x_{\alpha_{2}}x_{\alpha_{3}+\alpha_{4}}+\dfrac{3}{2}x_{\alpha_{4}}x_{\alpha_{2}+\alpha_{3}}-\dfrac{2}{3}x_{\alpha_{2}}x_{\alpha_{3}}x_{\alpha_{4}}$.
@@
Let us refer to Example Classification of coadjoint orbits for the maximal unipotent subgroup in the simple group of type $F_{4}$. There $D=\textup{supp}(\lambda)=\{\alpha_{4},\alpha_{3}+\alpha_{4},\alpha_{2},\alpha_{2}+2\alpha_{3}\}$. The sets $\zeta^{\prime}_{D,\alpha_{2}+2\alpha_{3}}$ and $\zeta^{\prime}_{D,\alpha_{2}+\alpha_{3}+\alpha_{4}}$ contain no nontrivial pairs. Indeed, there are no $\beta$ in $D$ such that $\beta>\alpha_{2}+2\alpha_{3}$ or $\beta>\alpha_{2}+\alpha_{3}+\alpha_{4}$. Since
$(\exp(\chi).\lambda)(e_{\alpha_{2}+\alpha_{3}})=2x_{\alpha_{3}}\lambda(e_{\alpha_{2}+2\alpha_{3}})=2x_{\alpha_{3}}(\exp(\chi).\lambda)(e_{\alpha_{2}+2\alpha_{3}})$

### v5

@@ 3 Classification of orbits
$\exp(x)\cdot\lambda=\sum_{\alpha\in\Phi^{+}}\lambda\left(\sum_{i=0}^{\infty}\dfrac{(-1)^{i}}{i!}\underbrace{[x,[x,\dots,[x,e_{\alpha}]\dots]]}_{\text{$i$ commutators}}\right)\otimes e_{\alpha}^{*}.$
Denote by $\chi$ the element $\chi=\sum_{\alpha\in\Phi^{+}}x_{\alpha}\otimes e_{\alpha}\in M$. For a form $\lambda\in\mathfrak{n}^{*}$, we will consider the element $\exp(\chi)\cdot\lambda\in M^{*}$. Let us also set $\overline{\mathfrak{n}}=\overline{\mathbb{F}}\otimes_{\mathbb{F}}\mathfrak{n}$, $\overline{N}=\overline{\mathbb{F}}\otimes_{\mathbb{F}}N$. Note that if $\widetilde{x}=\sum_{\alpha\in\Phi^{+}}\widetilde{x}_{\alpha}e_{\alpha}\in\overline{\mathfrak{n}}$ with $\widetilde{x}_{\alpha}\in\overline{\mathbb{F}}$, then $\exp(\widetilde{x})\cdot\lambda=(\exp(\chi)\cdot\lambda)(\widetilde{x}_{\alpha},\alpha\in\Phi^{+})$.
@@
Theorem 3.2.
Let $\lambda\in S$, $D=\operatorname{supp}(\lambda)$, and let $\gamma\in\Phi^{+}$ satisfy condition (1). Then there exists a rational function $F_{D,\gamma}\in\overline{\mathbb{F}}(y_{\beta},\beta\succ\gamma)$ such that
@@
where $\mu=\exp(\chi)\cdot\lambda\in M^{*}$.
Example $3.$
Consider the same support as in Example $2.$ Let
$D=\operatorname{supp}(\lambda)=\{\alpha_{4},\alpha_{3}+\alpha_{4},\alpha_{2},\alpha_{2}+2\alpha_{3}\}$.
@@
We will show that for this $\lambda$ and for all roots $\gamma$ satisfying condition (1), formula (2) holds. Let $\mu=\exp(\chi)\cdot\lambda,~\lambda=\sum_{\alpha\in\Phi^{+}}\lambda_{\alpha}e_{\alpha}^{*},~\mu=\sum_{\alpha\in\Phi^{+}}\mu_{\alpha}\otimes e_{\alpha}^{*}$. Let us write down explicitly the coordinates of interest. As in Example 2.7, we denote the structural constants in the relation $Ce_{\alpha+\beta}=[e_{\alpha},e_{\beta}]$ by $C_{\alpha,\beta}$. Their values can be found in [10]. We have
$\begin{split}\mu_{\alpha_{2}+2\alpha_{3}}&=\lambda(e_{\alpha_{2}+2\alpha_{3}})=\lambda_{\alpha_{2}+2\alpha_{3}};\\ \mu_{\alpha_{2}+\alpha_{3}+\alpha_{4}}&=0;\\ \mu_{\alpha_{2}+\alpha_{3}}&=-\lambda([x,e_{\alpha_{2}+\alpha_{3}}])=-\lambda([x_{\alpha_{3}}e_{\alpha_{3}},e_{\alpha_{2}+\alpha_{3}}])=-C_{\alpha_{3},\alpha_{2}+\alpha_{3}}x_{\alpha_{3}}\lambda(e_{\alpha_{2}+2\alpha_{3}})=\\ &=2x_{\alpha_{3}}\lambda_{\alpha_{2}+2\alpha_{3}};\\ \mu_{\alpha_{2}}&=\lambda(e_{\alpha_{2}})+\dfrac{1}{2}\lambda([x,[x,e_{\alpha_{2}}]])=\lambda_{\alpha_{2}}+\dfrac{1}{2}\lambda([x_{\alpha_{3}}e_{\alpha_{3}},[x_{\alpha_{3}}e_{\alpha_{3}},e_{\alpha_{2}}]])=\\ &=\lambda_{\alpha_{2}}+\dfrac{1}{2}C_{\alpha_{3},\alpha_{2}+\alpha_{3}}C_{\alpha_{3},\alpha_{2}}x_{\alpha_{3}}^{2}\lambda(e_{\alpha_{2}+2\alpha_{3}})=\lambda_{\alpha_{2}}-x_{\alpha_{3}}^{2}\lambda_{\alpha_{2}+2\alpha_{3}}=\\ &=\lambda_{\alpha_{2}}-\dfrac{\mu_{\alpha_{2}+2\alpha_{3}}^{2}}{4\mu_{\alpha_{2}+2\alpha_{3}}};\\ \mu_{\alpha_{3}+\alpha_{4}}&=\lambda(e_{\alpha_{3}+\alpha_{4}})=\lambda_{\alpha_{3}+\alpha_{4}};\\ \mu_{\alpha_{4}}&=\lambda(e_{\alpha_{4}})-\lambda([x,e_{\alpha_{3}+\alpha_{4}}])=\lambda_{\alpha_{4}}-\lambda([x_{\alpha_{3}}e_{\alpha_{3}},e_{\alpha_{3}+\alpha_{4}}])=\\ &=\lambda_{\alpha_{4}}-C_{\alpha_{3},\alpha_{4}}x_{\alpha_{3}}\lambda(e_{\alpha_{3}+\alpha_{4}})=\lambda_{\alpha_{4}}+x_{\alpha_{3}}\lambda_{\alpha_{3}+\alpha_{4}}=\lambda_{\alpha_{4}}+\dfrac{\mu_{\alpha_{2}+\alpha_{3}}\mu_{\alpha_{3}+\alpha_{4}}}{2\mu_{\alpha_{2}+2\alpha_{3}}}.\end{split}$
@@
In fact, the validity of formula (2) in this example could have been understood without writing out the coordinates explicitly. It is enough to examine Table 1. Since $\mu(e_{\alpha_{2}+\alpha_{3}})$ involves only $x_{\alpha_{3}}$ and $\lambda(e_{\alpha_{2}+2\alpha_{3}})$, and since $\lambda(e_{\alpha_{2}+2\alpha_{3}})=\mu(e_{\alpha_{2}+2\alpha_{3}})$, the variable $x_{\alpha_{3}}$ can be expressed as a rational function of $\mu(e_{\alpha_{2}+\alpha_{3}})$ and $\mu(e_{\alpha_{2}+2\alpha_{3}})$. Moreover, since $\mu(e_{\alpha_{2}})-\lambda(e_{\alpha_{2}})$ depends only on $x_{\alpha_{3}}$ and $\lambda(e_{\alpha_{2}+2\alpha_{3}})$, it can be expressed as a rational function of $\mu(e_{\alpha_{2}+\alpha_{3}})$ and $\mu(e_{\alpha_{2}+2\alpha_{3}})$. Furthermore, $\mu(e_{\alpha_{3}+\alpha_{4}})=\lambda(e_{\alpha_{3}+\alpha_{4}})$. As $\mu(e_{\alpha_{4}})-\lambda(e_{\alpha_{4}})$ involves only $x_{\alpha_{3}}$ and $\lambda(e_{\alpha_{3}+\alpha_{4}})$, it can be expressed as a rational function of $\mu(e_{\alpha_{2}+\alpha_{3}})$ and $\mu(e_{\alpha_{3}+\alpha_{4}})$.
@@
Now let us proceed to the proof of Theorem 3.2. First, we prove it in the special case of forms whose support consists of a single root. Any such form obviously belongs to $S$.
Lemma 3.4.
Theorem 3.2 holds for any $\lambda\in\mathfrak{n}^{*}$ with $|\operatorname{supp}(\lambda)|=1$.
Proof 3.5.
@@
Denote $D=\operatorname{supp}(\lambda)=\{\delta\}$. Note that in this case, condition (1) depends only on $D$, not on the specific $\lambda$, and it holds for those $\gamma$ that are not singular to $\delta$. Let $\mu=\exp(\chi)\cdot\lambda$. We need to prove that for any $\gamma\notin S_{\delta}$, the coordinate $\mu(e_{\gamma})$ is a function of the coordinates $\mu(e_{\beta}),\beta\succ\gamma$}. Observe that if $\gamma\succ\delta$, then $\mu(e_{\gamma})=0$, and if $\gamma=\delta$, then $\mu(e_{\delta})=\lambda(e_{\delta})$. For these two cases, the condition is obviously satisfied. It remains to prove it for $\gamma\prec\delta$.
@@
The coordinate $\mu(e_{\gamma})$ is a polynomial in the $x_{\beta}$ with coefficients from $\overline{\mathbb{F}}$. It does not contain the variable $x_{\beta_{2k-i+1}}$ if $\gamma\succ\beta_{i}$. Let $\beta_{2k-j+1}$ be the largest root in $S_{\delta}$ such that $x_{\beta_{2k-j+1}}$ appears in $\mu(e_{\gamma})$. Using formula (3), we express $x_{\beta_{2k-j+1}}$ in terms of $\mu(e_{\beta_{j}})$ and $x_{\alpha},\alpha<\beta_{2k-j+1}$}. Thus, $\mu(e_{\gamma})$ becomes a polynomial that no longer involves $x_{\beta_{2k-j+1}}$. We perform a similar elimination successively for $\beta_{2k-j+2},\beta_{2k-j+3}$, and so on, until $\beta_{2k}$. Note that at each step, a variable $x_{\beta_{i}}$ that was initially absent or already removed cannot reappear, because we proceed in decreasing order of the singular roots. As a result, $\mu(e_{\gamma})$ is represented as a polynomial that does not contain $x_{\beta},\beta\in S_{\delta}$, i.e.,
$\mu(e_{\gamma})=G(\mu(e_{\beta_{i}}),i=1,\ldots,j;~x_{\alpha},\alpha\notin S_{\delta})$
for some polynomial $G\in\overline{\mathbb{F}}[t_{\beta_{i}},i=1,\ldots,j;~x_{\alpha},\alpha\notin S_{\delta}]$. Observe that $\beta_{i}\succ\gamma$ for all $i=1,\ldots,j$.
@@
By our assumption, $\mu(e_{\gamma})$ cannot be expressed solely in terms of the subsequent coordinates. This implies that $G$ is not identically constant and actually depends on at least one variable $x_{\alpha}$. Then there exists a non-empty Zariski open subset $U\subseteq\overline{\mathbb{F}}^{j}$ such that for any $c=(c_{1},c_{2},\ldots,c_{j})\in U$, the specialized polynomial $G(c_{i},i=1,\ldots,j;~x_{\alpha},\alpha\notin S_{\delta})$ is non-constant. Fix such a tuple $c_{1},c_{2},\ldots,c_{j}\in\overline{\mathbb{F}}$. Choose an index $\varepsilon\notin S_{\delta}$ and constants $c_{\alpha}\in\overline{\mathbb{F}},\alpha\notin S_{\delta},\alpha\neq\varepsilon$, so that the one-variable polynomial
$g(x_{\varepsilon})=G(c_{i},i=1,\ldots,j;~x_{\varepsilon},c_{\alpha},\alpha\notin S_{\delta},\alpha\neq\varepsilon)$
is non-constant. The field is algebraically closed, for any chosen $c_{0}\in\overline{\mathbb{F}}$ there exists a $c_{\varepsilon}\in\overline{\mathbb{F}}$ such that $g(c_{\varepsilon})=c_{0}$. Fix $c_{0}$ and the corresponding $c_{\varepsilon}$. Now, choose arbitrary values $c_{j+1},\ldots,c_{2k}\in\overline{\mathbb{F}}$. Substituting the values $c_{\varepsilon},c_{\alpha},\alpha\notin S_{\delta}$ into formula (3), we can determine values for $x_{\beta_{2k}},x_{\beta_{2k-1}},\ldots,x_{\beta_{1}}$ sequentially, in such a way that $\mu(e_{\beta_{1}})=c_{1}$, $\mu(e_{\beta_{2}})=c_{2}$, $\ldots$ and $\mu(e_{\beta_{2k}})=c_{2k}$. For this assignment, we obtain
$\mu(e_{\gamma})=G(c_{i},i=1,\ldots,j;~c_{\varepsilon},c_{\alpha},\alpha\notin S_{\delta})=c_{0}.$
@@
Therefore, there exists a non-empty Zariski open subset $V\subseteq\overline{\mathbb{F}}^{2k+1}$ such that for any $(c_{0},c_{1},\ldots,c_{2k})\in V$, we can find an element $\overline{\mu}\in\overline{N}\cdot\lambda$ such that $\overline{\mu}(e_{\gamma})=c_{1}$, $\overline{\mu}(e_{\beta_{1}})=c_{2}$, $\ldots$, $\overline{\mu}(e_{\beta_{2k}})c_{2k}$. The orbit $\overline{N}\cdot\lambda$ is an algebraic variety. Our construction yields a surjection from $\overline{N}\cdot\lambda$ to $\overline{\mathbb{F}}^{2k+1}$, which implies $\dim\overline{N}\cdot\lambda\geq 2k+1$. However,
$\dim\overline{N}\cdot\lambda=\dim\overline{\mathfrak{n}}\cdot\lambda=\dim\mathfrak{n}\cdot\lambda=|S_{\delta}|=2k.$
This contradiction completes the proof.
@@
We introduce further notation. Let $\alpha,\beta\in\Phi^{+}$. Denote $T_{\alpha,\beta}(x):=(\exp(\chi).e_{\beta}^{*})(e_{\alpha})$ in $\overline{\mathbb{F}}[x_{\gamma},\gamma\in\Phi^{+}]$. It is clear that if $\lambda\in\mathfrak{n}^{*},\alpha\in\Phi^{+}$, then
@@
$(\exp(\chi)\cdot\lambda)(e_{\alpha})$ $=\left(\exp(\chi)\cdot\left(\sum_{\beta\in\operatorname{supp}(\lambda)}\lambda(e_{\beta})e_{\beta}^{*}\right)\right)(e_{\alpha})$
$=\sum_{\beta\in\operatorname{supp}(\lambda)}\lambda(e_{\beta})(\exp(\chi).e_{\beta}^{*})(e_{\alpha})$
$=\sum_{\beta\in\operatorname{supp}(\lambda)}\lambda(e_{\beta})T_{\alpha,\beta}(x).$
@@
Now consider the following construction (cf. the construction of the set $\zeta_{D,\gamma}$ used in the proof of the sufficient condition in Proposition 2.3). Let $\lambda\in S,~D=\operatorname{supp}(\lambda)$, and let the root $\gamma\in\Phi^{+}$ satisfy condition (1). Starting from $D$ and $\gamma$, we construct a set $\zeta^{\prime}_{D,\gamma}\subset\Phi^{+}\times\Phi^{+}$: by definition, it consists of those pairs $(\alpha,\beta)$ for which the polynomial $T_{\alpha,\beta}(x)$ can be expressed as a rational function of $\{(\exp(\chi)\cdot\lambda)(e_{\delta})\}_{\delta\succ\gamma}$.
Example 3.6.
@@
Table 1 helps visualize the structure of the polynomials $T_{\alpha,\beta}(x)$. One simply examines the cell in row $\alpha$ and column $\beta$. For each pair $\gamma,\delta\in\Phi^{+}$ such that $\gamma+\delta\in\Phi^{+}$, denote the corresponding non-zero structure constant by $C_{\gamma,\delta}$, so that $e_{\gamma},e_{\delta}]=C_{\gamma,\delta}e_{\gamma+\delta}$. Each monomial in $T_{\alpha,\beta}(x)$ has the form
@@
where $\gamma_{1},\ldots,\gamma_{n}\in\Phi^{+},~\alpha+\gamma_{1}+\ldots+\gamma_{i}\in\Phi^{+}$ for all $i$ from 11 to $n$, and $\alpha+\gamma_{1}+\ldots+\gamma_{n}=\beta$. Hence, it suffices to consider all possible decompositions of $\beta-\alpha$ into a sum of positive roots. For example,
$T_{\alpha_{3},\alpha_{2}+2\alpha_{3}+\alpha_{4}}(x)=-x_{\alpha_{2}+\alpha_{3}+\alpha_{4}}+\dfrac{1}{2}x_{\alpha_{2}}x_{\alpha_{3}+\alpha_{4}}+\dfrac{3}{2}x_{\alpha_{4}}x_{\alpha_{2}+\alpha_{3}}-\dfrac{2}{3}x_{\alpha_{2}}x_{\alpha_{3}}x_{\alpha_{4}}$.
@@
Let us refer to Example $3.$ There $D=\operatorname{supp}(\lambda)=\{\alpha_{4},\alpha_{3}+\alpha_{4},\alpha_{2},\alpha_{2}+2\alpha_{3}\}$. The sets $\zeta^{\prime}_{D,\alpha_{2}+2\alpha_{3}}$ and $\zeta^{\prime}_{D,\alpha_{2}+\alpha_{3}+\alpha_{4}}$ contain no nontrivial pairs. Indeed, there are no $\beta$ in $D$ such that $\beta>\alpha_{2}+2\alpha_{3}$ or $\beta>\alpha_{2}+\alpha_{3}+\alpha_{4}$. Since
$(\exp(\chi)\cdot\lambda)(e_{\alpha_{2}+\alpha_{3}})=2x_{\alpha_{3}}\lambda(e_{\alpha_{2}+2\alpha_{3}})=2x_{\alpha_{3}}(\exp(\chi)\cdot\lambda)(e_{\alpha_{2}+2\alpha_{3}})$